OpenStudy (anonymous):

If F(x) = x² + 5x and G(x) = 2x + 1, find F(5) + G(6).

OpenStudy (anonymous):

make F(5) first

OpenStudy (anonymous):

F(5) is F(x)=x^2 + 5x if you put x=5

OpenStudy (anonymous):

so its F(5)=5^2 +5*5

OpenStudy (anonymous):

everything ok till now?

OpenStudy (anonymous):

yes thank u

OpenStudy (anonymous):

i still dont get t

OpenStudy (anonymous):

how much is F(5)?

OpenStudy (anonymous):

did you calculate it?